OpenStudy (anonymous):

For any of you calc II fans: Im stuck on this! sum_{n=3}^{infty} (n^3)/(2n+1)! I believe this converges, but using a ratio test, it appears to be infinity over infinity, which is inconclusive. Im doing something wrong. Please help me out someone!

ganeshie8 (ganeshie8):

what do u get after simplifying the ratio ?

OpenStudy (anonymous):

I get 2n+1/2n+2, which applying a lim as x goes to infinity would be inconclusive, infinity over infinity.

ganeshie8 (ganeshie8):

\[\rm \begin{align} \lim\limits_{n\to \infty} \left|\dfrac{a_{n+1}}{a_n}\right| & = \lim\limits_{n\to \infty}\left|\dfrac{(n+1)^3/(2(n+1)+1)!}{n^3/(2n+1)!}\right| \\~\\ & = \lim\limits_{n\to \infty}\left|\dfrac{(n+1)^3/(2n+3)!}{n^3/(2n+1)!}\right| \\~\\ \end{align}\]

OpenStudy (anonymous):

Can I pull out n! from (2n+1)! or would it be 2n! ? I am not very familiar with the rules of factorials.

ganeshie8 (ganeshie8):

\[n! = n\cdot (n-1)!\]

ganeshie8 (ganeshie8):

\[n! = n\cdot (n-1)! = n\cdot(n-1)\cdot (n-2)!\]

ganeshie8 (ganeshie8):

you can keep going like that till you reach 1

ganeshie8 (ganeshie8):

lets rearrange the fractions first \[\rm \begin{align} \lim\limits_{n\to \infty} \left|\dfrac{a_{n+1}}{a_n}\right| & = \lim\limits_{n\to \infty}\left|\dfrac{(n+1)^3/(2(n+1)+1)!}{n^3/(2n+1)!}\right| \\~\\ & = \lim\limits_{n\to \infty}\left|\dfrac{(n+1)^3/(2n+3)!}{n^3/(2n+1)!}\right| \\~\\ & = \lim\limits_{n\to \infty}\left|\dfrac{(n+1)^3(2n+1)!}{n^3(2n+3)!}\right| \\~\\ \end{align}\]

OpenStudy (anonymous):

I just tried this on paper, and I end up with \[(2n+1)!/(2n+3)!\]

OpenStudy (anonymous):

can i pull out a factorial n! from these?

ganeshie8 (ganeshie8):

you cannot cancel (n+1)^3 and n^3

OpenStudy (anonymous):

I still do not understand how to get anywhere with this. I think you have the correct fraction, I can arrive at that as well, however I don't see how to obtain the ratio from it.

ganeshie8 (ganeshie8):

lets split (2n+3)! as (2n+3)(2n+2)(2n+1)! \[\rm \begin{align} \lim\limits_{n\to \infty} \left|\dfrac{a_{n+1}}{a_n}\right| & = \lim\limits_{n\to \infty}\left|\dfrac{(n+1)^3/(2(n+1)+1)!}{n^3/(2n+1)!}\right| \\~\\ & = \lim\limits_{n\to \infty}\left|\dfrac{(n+1)^3/(2n+3)!}{n^3/(2n+1)!}\right| \\~\\ & = \lim\limits_{n\to \infty}\left|\dfrac{(n+1)^3(2n+1)!}{n^3\color{red}{(2n+3)!}}\right| \\~\\ & = \lim\limits_{n\to \infty}\left|\dfrac{(n+1)^3(2n+1)!}{n^3\color{Red}{(2n+3)(2n+2)(2n+1)!}}\right| \\~\\ \end{align}\]

ganeshie8 (ganeshie8):

cancel the (2n+1)! top and bottom

OpenStudy (anonymous):

ok, now I am at \[(n+1)^3/(n^3*(2n+3)(2n+2))\]

ganeshie8 (ganeshie8):

yes take the limit

OpenStudy (anonymous):

this is still infinity over infinity

ganeshie8 (ganeshie8):

notice that degree of numerator = 3 degree of denominator = 4

ganeshie8 (ganeshie8):

the higher degree denominator clearly pulls the value of fraction to 0 as n -> infinity

OpenStudy (anonymous):

I get 5 as the degree in the denominator. I foiled the (2n+3)(2n+2) which is 4n^2+10n+6 and distributed n^3 to have n^5 as the highest degree. Am I doing this wrong as well?

ganeshie8 (ganeshie8):

Oh right, degree of denominator is 5

OpenStudy (anonymous):

ok, I know this doesn't change it. So, the faster approach of the denominator makes this become zero.

OpenStudy (anonymous):

Thank you for your help ganeshie8